Determine values for all resistors in the biasing circuit of a Common Source n- MOSFET Amplifier configuration. VDD VDD 9V 9VAlso please solve for the Gain
Answer #1

2 Given - 10 Im mosfet q, is in saturation Io = kn ( Vas Ven? deimd = mm. (Vas - Yto) 1 = (vas - Venda NT Ima 2.ma 26mV 9m 0.Rot Rs = NDO OS 9-3 = 6ks 2 m [RO = 4.7562 circuit co - from Vo = -9m Vgs Ro. Vin = Vgs + Om Vgs Rs. AU -9m. Ro. Hogs (+9m Ri